Paul wants to invest a total of $5000 in two accounts, one paying 4% annual simple interest and the other 3%. If he wants to earn $165 annua; interest, how much should he invest at each rate?
Answer by checkley71(8403)

.04X+.03(5000-X)=165
.04X+150-.03X=165
.01X=165-150
.01X=15
X=15/.01
X=$1,500 SHOULD BE INVESTED @ 4%.
5000-1500=$3,500 INVESTED @ 3%.
PROOF
.04*1500+.03*3500=165
60+105=165
165=165